OpenStudy (qaisersatti):

Homework 9 Pt 1 (ii,iii & v) Pt 2 (iv,v,vii) any help please

OpenStudy (nurali):

Question 1 2. 70 3. 70 5. -219.9 Question 2 didn't Complete please tell me answer of question 2 i,ii,iii,vi

OpenStudy (nurali):

if correct please click best response

OpenStudy (anonymous):

H9P1 0.250122093132 7.5 7.5 0 -23.5619449019 0 0 0 H9P2 L*C L/R 1 HELP Help (1/R)*sqrt(L/C) HELP H9P3 b/(2*m) 1/(sqrt(m*1/k)) 1/(2*pi*sqrt(m*1/k)) (2*m)/(sqrt(m*(1/k)))

OpenStudy (anonymous):

Reviwe H9P1 a) 0.250122093132 b) 7.5 c) 7.5 d) 0 e) -23.5619449019 f) 0 g) 0 h) 0 H9P2 a) L*C b) L/R c) 1 d) sqrt(1/(L*C))/(2*pi) e) 1/(2*R*C) f) (1/R)*sqrt(L/C) g) 2*sqrt(L/C) H9P3 a) b/(2*m) b) 1/(sqrt(m*1/k)) c) 1/(2*pi*sqrt(m*1/k)) d) (2*m)/(sqrt(m*(1/k)))

OpenStudy (anonymous):

At the time just before the impulse happens what is the current iL(5.0−), in Amperes, through the inductor? Equation???

OpenStudy (anonymous):

iL(t) = A*cos(2*pi*f*t) + B*sin(2*pi*f*t) vC(t) = L*diL/dt = L*2*pi*f*(-A*sin(2*pi*f*t) + B * cos(2*pi*f*t)) The initial conditions mean that A = 1, and B = 0, (in my case) So, iL (t) = cos(2*pi*f*t), vC (t) = -L*2*pi*f*sin(2*pi*f*t) iL(5) = cos(2*pi*0.25*5) = cos(2.5*pi) = 0. If you have another values, change only.

OpenStudy (qaisersatti):

My Q No 1 (ii,iii & v) is not coming correct can anyone comment. The current source puts out an impulse of area A=2/π=0.64 Coulombs at time t=5.0s. At t=0 the state is: vC(0)=0.0 and iL(0)=1.0. The equation governing the evolution of the inductor current in this circuit is d2iL(t)dt2+1LCiL(t)=ALCδ(t−5.0) (ii) At the initial time what is the total energy, in Joules, stored in the circuit? (iii) At the time just before the impulse happens t=5.0− what is the total energy, in Joules, stored in the circuit? (v) At the time just before the impulse happens what is the voltage vC(5.0−), in Volts, across the capacitor? please comment...

OpenStudy (anonymous):

Thats the same values and questions in my, i don´t know what happens.

OpenStudy (anonymous):

try this energy stored=½*L*iL^2 (b and c) vC(t) = -L*2*pi*f*sin(2*pi*f*t) for (e)

OpenStudy (qaisersatti):

@Jose.oliveira2 thanks brother its already done...thanks